# DyFFPAD: Dynamic Fusion of Convolutional and Handcrafted Features for Fingerprint Presentation Attack Detection

### v2

@@ Abstract
Automatic fingerprint recognition systems suffer from the threat of presentation attacks due to their wide range of deployment in areas including national borders and commercial applications. A presentation attack can be performed by creating a spoof of a user’s fingerprint with or without their consent. This paper presents a dynamic ensemble of deep CNN and handcrafted features to detect presentation attacks in known-material and unknown-material protocols of the livness detection competition. The proposed presentation attack detection model, in this way, utilizes the capabilities of both deep CNN and handcrafted features techniques and exhibits better performance than their individual performances. The proposed method is validated using benchmark databases from the Liveness Detection Competition in 2015, 2017, and 2019, yielding overall accuracy of 96.10%, 96.49%, and 94.99% on them, respectively. The proposed method outperforms state-of-the-art methods in terms of classification accuracy.
Index Terms:
Presentation Attack Detection, Hybrid Architecture, Handcrafted Features, Deep CNN, Fingerprint Biometrics.
@@ 1 Introduction
Automatic Fingerprint Recognition Systems (AFRS) are most widely used for the purpose of person authentication and verification. Their user-friendliness and cost-effectiveness make them popular while validating the identity of persons at airports, national borders, and distribution of government-funded aid. The utilization of these systems in a wide range of applications, also makes them vulnerable to some internal and external security threats. Presentation Attack (PA), or spoofing, is an external attack where an artificial artifact of a user’s finger is presented to the sensor of an Automated Fingerprint Recognition System (AFRS). PAs or spoofs can be created either by a cooperative method or a non-cooperative method of spoofing. Fabrication materials such as latex, woodglue, gelatine, silicon, ecoflex, etc. are available at a very low cost to create the spoof of a fingerprint. Fingerprint Presentation Attack Detection (FPAD) serves as a countermeasure to these attacks and to empower an AFRS against them. The recent FPAD methods suggested by various researchers are categorized into two broad classes including hardware-based and software-based methods. Hardware-based methods are quite expensive due to the utilization of additional sensing devices that measure natural properties such as heart rate, odor, temperature, etc. The utilization of these sensors makes hardware-based methods less user-friendly and quite expensive for an organization to use in access-control applications. On the other side, software-based methods require only fingerprint samples which makes them friendly to the user and cost-effective to the organization as compared to hardware-based methods. Current software-based methods can be further categorized into perspiration and pore based-methods [9], [2] statistical and handcrafted feature-based methods [32], [31], [37], [3] and deep-learning based methods [5], [4], [28]. Perspiration-based methods suffer from dryness in the environment which sometimes causes the rejection of a genuine fingerprint sample. Similarly, pore-based methods necessitate the sensing device to capture a high-definition image of the fingertip. This necessity impacts the overall cost of these methods. Existing works utilizing statistical and handcrafted feature extract some predefined features from the input fingerprint sample to classify them as live or spoof. These methods are affected by the quality of the fingerprint samples. However, some of these methods including Sharma et al.[32], Sharma et al. [31] have shown good FPAD capability while the spoof sample is created using known material but they have not been tested properly on the spoofs fabricated with unknown materials. Deep learning methods consist of a classifier that possesses convolutional layers as a feature extractor. These layers have an unmatched capability of extracting minute features from the input samples. In recent times, various methods based on deep learning have been suggested by various researchers which have exhibited better performance than other software-based methods but they also suffer while being tested with the spoofs fabricated with unknown materials.
@@
In our proposed work, a dynamic ensemble of handcrafted and deep features based model is presented. This Dynamic Fusion of convolutional and handcrafted features for Fingerprint Presentation Attack Detection (DyFFPAD), is an end-to-end approach that dynamically incorporates deep and handcrafted features. The model is build using two sub-models that work in collaboration to classify the live and spoof fingerprint samples. The first sub-model is a feature extractor model comprising of Deep Neural Network (DNN) which is empowered with an image descriptor namely Local Phase Quantization (LPQ) and a set of quality-based handcrafted features including Ridge Valley Clarity (RVC), Gabor quality, Frequency Domain Analysis (FDA), and Orientation Flow (OFL), etc. On the other hand, the second sub-model is a DenseNet architecture’s feature base attached to a custom-made DNN. DenseNet has shown remarkable performance while being tested on some well-known image classification problem databases including MNIST [7], CIFAR [36], and imagenet [6]. The output of both the sub-models is combined and fed to another DNN which has a single neuron in the last for the binary classification. The TSNE plots depicted in Fig. 1 show that the proposed method is able to distinguish between live samples, and spoof samples irrespective of their fabrication materials. The proposed model exhibits the same performance on different datasets and its efficacy does not affect whether the spoofs are fabricated with known or unknown materials. The performance of the proposed method is validated using Liveness Detection Competition (LivDet) databases, demonstrating superior results compared to state-of-the-art methods in both intra-sensor same-material and intra-sensor cross-material protocols. A detailed comparison with the state-of-the-art methods indicates the supremacy of the proposed method while detecting the PAs in benchmark protocols of FPAD.
The main contributions of this paper are highlighted below.
@@
2. A dynamic fusion of handcrafted and convolutional features is proposed which addresses the limitations that occur in static fusion.
@@
5. From TSNE representations, it is evident that the proposed model demonstrates consistency while detecting the spoofs fabricated with known and unknown materials.
The remainder of this paper is organized as follows: Section 2 reviews state-of-the-art methods proposed by researchers for the detection of PAs, highlighting their advantages and limitations. Section 3 details the design and functionality of the proposed architecture. Section 5 presents the experimental results, followed by a comparative analysis. Finally, the conclusion is provided in Section 6.
@@ 2 Related Work
PAs represent the most significant security threat to the AFRS The applications of AFRS in security and commercial areas brought the attention of researchers toward the development of cost-effective and user-friendly solutions to this problem. In this section, methods put forth by researchers to keep the AFRS safe against PAs are discussed. Existing methods utilizes perspiration and pore features, statistical and handcrafted features, and deep learning features. The methods that fall into these categories and their limitations are discussed in this section.
@@ 2.1 Perspiration and pore based-methods
Perspiration is a natural property in human skin that is caused by the presence of tiny holes or pores. As spoofs created with different materials do not have this natural property, it could be utilized to develop a method that distinguishes between live and spoof fingerprints based on this property. In this attempt, Derakshini et al. [8] leveraged the presence of perspiration and sweat diffusion patterns to differentiate between live and spoofed fingerprints. Further, Abhyankar et al. [2] proposed a wavelet-based approach to detect PAs using the fingerprint’s perspiration property. A custom-made fingerprint database is utilized to validate the performance of the proposed method. The count of pores within a live fingerprint and its counterfeits crafted from various materials can exhibit variability because pores are hard to reflect in the spoofs during the fabrication process. Espinoza [9] utilized this dissimilarity as a distinguishing feature to detect PAs. The proposed method is validated on a custom-made dataset. Marcialis et al. [21] also proposed a similar method that detects pores from multiple fingerprints captured at an interval of five seconds. The number of pores is further in the subsequent samples is further utilized as a feature to detect PAs. The proposed method is validated on a custom-made database having 8,960 live and 7,760 spoof samples. Manivanan et al. [20] utilized high-pass filtering followed by correlation filtering techniques [22] to identify active sweat pores on fingerprint ridges. The results demonstrate distinguishable correlation peaks corresponding to the sweat pores on the ridges. Regardless of how the perspiration pattern is employed to detect PAs, its presence is dependent on the temperature of the surrounding environment. Sometimes, even a live finger does not exhibit this property in a dry environment which occasionally prompts the FPAD method to mis-classify this live sample as a spoof. Similarly, the pore-based methods cost high due to the utilization of a high-resolution sensing device that is capable of collecting fingerprint samples with higher dpi (>=>=> =1000 pixels per inch). These factors collectively make perspiration-based and pore-based methods less economical and less convenient to use.
2.2 Statistical and handcrafted feature based-methods
@@
The inherent disparities between the natural attributes of a finger’s skin and counterfeit spoofs, such as variations in color and moisture levels, manifest in the quality of collected samples. This phenomenon prompted researchers to utilize the quality of captured fingerprint samples as a distinguishing feature. This section describes some of the approaches that fall within this category. Park et al. [27] employed a range of statistical features, encompassing deviation, hyper-flatness, average brightness, standard deviation, variance, skewness, hyper-skewness, kurtosis and differential image features, for the training of a Support Vector Machine (SVM) classifier, enabling detection of live and spoof samples. The efficacy of the proposed approach is evaluated through experimentation on the ATVSFFp database, comprising 272 genuine and 270 spoof fingerprint samples. In a similar attempt, Xia et al. [37] proposed an image descriptor that captures both intensity variance and gradient properties to create a combined feature vector. This feature vector is then used to train an SVM classifier for spoof detection. The proposed method is validated using the LivDet 2011, 2013, and 2015 databases. Kim et al. [18] proposed an image descriptor that utilizes the local coherence of fingerprints as a feature to train an SVM classifier. LivDet 2009, 2011, 2013, and 2015 and ATVSFFp databases are utilized by the authors to validate their work. Yuan et al. [40] introduced a novel approach that utilizes gradient properties for image analysis. This method involves generating image gradient matrices through diverse quantization operators which are derived from the computation of two co-occurrence matrices using the Laplacian operator. These computed matrices are then employed to train a back-propagation neural network. The effectiveness of the proposed method is demonstrated using the LivDet 2013 database. Gragnaniello et al. [13] employed the Weber Local Descriptor (WLD) to extract digital excitation and gradient information from fingerprint samples. These extracted features are subsequently used to train an SVM classifier. The proposed method is validated using the LivDet 2009 and 2011 databases. In another work, Gragnaniello et al. [14] proposed a novel image descriptor named Local Contrast Phase Descriptor (LCPD) which is based on a component of spatial domain and rotation invariant version of Local Phase Quantization (LPQ). The proposed method is validated on LivDet 2011 database. Rattani et al. [29] proposed an incremental learning based approach that consist of a novel material detector along with the classifier. This detector classifies input fingerprint samples into three categories: live, spoof, and unknown. Samples classified as ’novel’ are further used to train the classifier. The proposed method is validated using LivDet 2011 database. In continuation to their previous work [29], Rattani et al. [30] proposed a novel approach that employs a Weibull-calibrated SVM for the classification of fingerprint samples. This approach integrates one-class and binary SVMs, demonstrating a significant improvement over their previous method [29]. Since the finger skin and spoof fabrication materials exhibit different levels of elasticity the resultant fingerprints have some differences in terms of quality. The difference in the live and spoof samples fabricated with different materials results in uneven ridge and valley widths and the fabrication process introduces other quality irregularities in the spoofs. In an attempt to develop a quality-based FPAD model, Sharma et al. [32] utilized some fingerprint quality features for the detection of PAs. They extracted Ridge and Valley Smoothness (RVS), Ridge and Valley Clarity (RVC), Frequency Domain Analysis (FDA), and Orientation Certainty Level (OCL) which are combined to form a feature vector. The feature vector is subsequently utilized to train a Random-Forest (RF) classifier. The proposed method is validated on a variety of databases including LivDet 2009, 2011, 2013, and 2015. Similarly, Sharma et al. [31] proposed a novel variant of the Local Binary Pattern (LBP) known as the Local Adaptive Binary Pattern (LABP). This feature is combined with the existing Complete Local Binary Pattern (CLBP) and Binary Statistical Image Features (BSIF) to train both an SVM classifier and a Deep Neural Network (DNN). The effectiveness of the proposed method is validated across multiple databases, including LivDet 2009, 2011, 2013, and 2015. The quality of the fingerprint samples plays a vital role in PA detection by these methods. Some of the methods including [32], [31] [18] exhibit good performance while being evaluated in the known material scenario but have not been properly tested in the cross-material protocol of FPAD.
@@ 2.3 Deep learning-based-methods
Deep Convolutional Neural Networks (CNNs) excel at extracting minute features from input image samples. These models have shown remarkable classification performance on image datasets such as MNIST [7], CIFAR [36], and ImageNet [6]. This ability of the CNNs has prompted researchers to explore their application in the detection of PAs as well. The state-of-the-art methods that use deep CNNs for FPAD are discussed in this section. Park et al. [26] extracted multiple patches from segmented fingerprints and trained a custom-made CNN composed of three convolutional and one fully connected layer. The proposed model utilizes a voting strategy instead of score-level fusion to compute the overall liveness score for an input fingerprint sample. The proposed method is validated using LivDet 2009 database. In a similar attempt, Toosi et al. [34] applied fingerprint foreground segmentation on fingerprint images followed by patch extraction. The extracted patches are further used to train a pre-trained Alexnet [19] CNN classifier. The proposed method is validated using LivDet 2011 and 2013 databases. Nogueira et al. [10] trained two separate SVMs using the features extracted with Local Binary Pattern (LBP) image descriptor and a single-layer CNN respectively. They also incorporated various data augmentation techniques including image rotation and scaling. The limitation of this work is the feature extraction and the classification tasks are performed as separate tasks which do not allow the entire system to learn their parameters together. In another work, Nogueira et al. [24] utilized pre-trained CNN classifiers in their work. They employed VGG, Alexnet, CNN, along with an SVC trained with well-known handcrafted features and applied score-level fusion for the detection of PAs. The proposed method is validated using Livdet 2009, 2011, and 2013 databases. Uliyan et al. [35] proposed a deep feature-based approach for detecting PAs. Their method utilizes a Deep Boltzmann Machine (DBM) for feature extraction and a Restricted Boltzmann Machine (RBM) to establish relationships between these features. Yuan et al. [39] proposed a multi-modal approach utilizing multiple pre-trained CNNs to detect the PAs. The proposed method is validated using LivDet 2011, 2013, 2015, and NUAA face databases. Chugh et al.[5] proposed a method that utilizes minutiae-centered patches for training and testing on the MobileNet classifier. The global liveness score for an input sample is computed by the fusion of the liveness score generated by the classifier for all the patches. The proposed method is validated using LivDet 2011, 2013, 2015, and Michigan State University’s (MSU) FPAD database. Zhang et al. [42] proposed a novel CNN architecture with a series of improved residual connected blocks. It reduces the problem of over-fitting and also takes less processing time. The proposed method is validated on Livdet 2013 and 2015 databases. Since, novel fabrication materials are being discovered nowadays, generalizing an FPAD model to detect PAs in an open-set paradigm is a challenging issue. Chugh et al. [4], in continuation to their previous work [5], proposed an approach for detecting spoofs fabricated with unknown materials. They have presented an image synthesis technique for generating new fingerprint patches which augments the training of MobileNet classifier representing the unknown materials spoofs while experimenting with LivDet 2017, ATVSFFp, and MSU-FPAD databases. Generative Adversarial Networks can also be utilized to augment the training database with spoofs fabricated with unknown materials. In a similar attempt, Rai et al. in [28] utilized Wasserstein GAN to generate fingerprint patches for better training of the DenseNet classifier against the spoofs fabricated with unknown material. The proposed method is validated on LivDet 2015, 2017, and 2019 databases. Although the Deep learning-based methods are quite effective for image classification problems still, they suffer from low classification accuracy in the area of FPAD while being tested in the cross-material protocol of FPAD. The detailed literature survey concludes that the deep learning-based methods are proven to be superior as compared with the methods belonging to other categories but their performance is required to be improved in cross-material protocol of FPAD. This objective can be attained by developing a hybrid model that utilizes the power of both convolutional and handcrafted features. The proposed model should also be efficient enough to classify the input sample in a minimum amount of time to be incorporated with the AFRS.

### v5

@@ Abstract
Automatic fingerprint recognition systems suffer from the threat of presentation attacks due to their wide range of deployment in areas including national borders and commercial applications. A presentation attack can be performed by creating a spoof of a user’s fingerprint with or without their consent. This paper presents a deep neural network-based solution to safeguard fingerprint recognition systems against such attacks. The proposed dynamic ensemble of deep CNN and handcrafted features to detect presentation attacks in known-material and unknown-material scenarios of the liveness detection. The proposed presentation attack detection model, in this way, utilizes the capabilities of both deep CNN and handcrafted features techniques and exhibits better performance than their individual performances. We have validated our proposed method on benchmark databases from the Liveness Detection Competition in 2015, 2017, and 2019, yielding overall accuracy of 96.10%, 96.49%, and 94.99% on them, respectively. The proposed method outperforms state-of-the-art methods in terms of classification accuracy.
keywords:
[…]
{highlights}
This paper presents a Deep Neural Network-based method to protect the Automatic Fingerprint Recognition System (AFRS) from Presentation Attacks (PAs).
It presents a novel fusion of convolutional and handcrafted features which address the limitations that occur in the static fusion of these features.
The proposed method performs better than state-of-the-art methods in benchmark scenarios.
@@ 1 Introduction
Automatic Fingerprint Recognition Systems (AFRS) are most widely used for the purpose of person authentication and verification Nagaty (2001). Their user-friendliness and cost-effectiveness make them popular while validating the identity of persons at airports, national borders, and distribution of government-funded aid. The utilization of these systems in a wide range of applications, also makes them vulnerable to some internal and external security threats. Presentation Attack (PA), or spoofing, is an external attack where an artificial artifact of a user’s finger is presented to the sensor of an Automated Fingerprint Recognition System (AFRS). PAs or spoofs can be created either by a cooperative method or a non-cooperative method of spoofing. Fabrication materials such as latex, woodglue, gelatine, silicon, ecoflex, etc. are available at a very low cost to create the spoof of a fingerprint. Fingerprint Presentation Attack Detection (FPAD) serves as a countermeasure to these attacks and to empower an AFRS against them. The recent FPAD methods suggested by various researchers are categorized into two broad classes including hardware-based and software-based. The methods based on hardware are quite expensive due to the utilization of additional sensing devices that measure natural properties such as heart rate, odor, temperature, etc. The utilization of these sensors makes hardware-based methods less user-friendly and quite expensive for an organization to use in access-control applications. On the other side, software-based methods require only fingerprint samples which makes them friendly to the user and cost-effective to the organization as compared to hardware-based methods. Current software-based methods can be further categorized into perspiration and pore based-methods Espinoza and Champod (2011), Abhyankar and Schuckers (2009) statistical and handcrafted feature-based methods Sharma and Dey (2019), Sharma and Selwal (2021), Xia et al. (2020), Baishya et al. (2023) and deep-learning based methods Chugh and Jain (2021), Chugh et al. (2018), Rai et al. (2024), and Rai et al. (2025). Perspiration-based methods suffer from dryness in the environment which sometimes causes the rejection of a genuine fingerprint sample. Similarly, pore-based methods necessitate the sensing device to capture a high-definition image of the fingertip. This necessity impacts the overall cost of these methods. Existing works utilizing statistical and handcrafted feature extract some predefined features from the input fingerprint sample to classify them as live or spoof. These methods are affected by the quality of the fingerprint samples. However, some of these methods including Sharma and Dey (2019), Sharma and Selwal (2021) have shown good FPAD capability while the spoof sample is created using known material but they have not been tested properly on the spoofs fabricated with unknown materials. Deep learning methods consist of a classifier that possesses convolutional layers as a feature extractor. These layers have an unmatched capability of extracting minute features from the input samples. In recent times, various deep learning methods have been suggested by various researchers which have exhibited better performance than other software-based methods but they also suffer while being tested with the spoofs fabricated with unknown materials.
@@
In our proposed work, a dynamic ensemble of handcrafted and deep features-based model is presented. This Dynamic Fusion of convolutional and handcrafted features for Fingerprint Presentation Attack Detection (DyFFPAD), is an end-to-end approach that dynamically incorporates deep and handcrafted features. The model is build using two sub-models that work in collaboration to classify the live and spoof fingerprint samples. The first sub-model is a feature extractor model comprising a Deep Neural Network (DNN) which is empowered with an image descriptor namely Local Phase Quantization (LPQ) and a set of quality-based handcrafted features including Ridge Valley Clarity (RVC), Gabor quality, Frequency Domain Analysis (FDA), and Orientation Flow (OFL), etc. On the other hand, the second sub-model is a DenseNet architecture’s feature base attached to a custom-made DNN. DenseNet has shown remarkable performance while being tested on some well-known image classification problem databases including MNIST Deng (2012), CIFAR Wu et al. (2020), and imagenet Deng et al. (2009). The output of both the sub-models is combined and fed to another DNN which has a single neuron in the last for the binary classification. The proposed model addresses the issue of developing a dynamic fusion of convolutional and handcrafted features in a better way as it does not require any intermediary stage of data preparation and training of multiple classifiers, unlike other static fusion-based models presented in (Nguyen et al., 2018b, a; Yadav et al., 2018). The TSNE plots depicted in Fig. 1 show that the proposed method can distinguish between live samples, and spoof samples irrespective of their fabrication materials. The proposed model shows the same performance on different datasets and its efficacy is not affected by known or unknown materials used for spoof fabrication. The proposed method is evaluated using Liveness Detection Competition (LivDet) databases, demonstrating superior results compared to state-of-the-art methods for the evaluation protocol of intra-sensor same-material and intra-sensor cross-material. A detailed comparison with the state-of-the-art methods indicates the supremacy of the proposed method while detecting the PAs in benchmark protocols of FPAD.
The major contributions of our proposed work are highlighted below.
@@
2. A dynamic fusion of handcrafted and convolutional features is proposed which addresses the limitations that occur in static fusion of these features.
@@
5. From TSNE representations, it is evident that the proposed model demonstrates consistency while detecting spoofs fabricated with known and unknown materials.
The rest of this paper is arranged in following order: Section 2 presents related work by researchers for the detection of PAs, highlighting their advantages and limitations. Section 3 provides the details of architecture and design of proposed method. Section 4 presents the experimental setup and protocol details while Section 5 provides the experimental results, followed by a comparative analysis. Finally, the conclusion is provided in Section 6.
@@ 2 Related Work
PAs poses a significant security threat to the AFRS which has brought the attention of researchers toward the development of cost-effective and user-friendly solutions to this problem. We have provided an overview of some of the prominent methods put forth by researchers to keep the AFRS safe against PAs in this section. Existing methods utilizes perspiration and pore features, statistical and handcrafted features, and deep learning features. The methods that fall into these categories and their limitations are discussed in this section.
@@ 2.1 Perspiration and pore based-methods
Perspiration is a natural phenomenon in human skin that is caused by the presence of tiny holes or pores. As spoofs created with different materials do not have this natural property, it could be utilized to develop a method that distinguishes between live and spoof fingerprints based on this property. In this attempt, Derakhshani et al. (2003) leveraged the presence of perspiration and diffusion patterns of sweat to differentiate between live and spoof fingerprints. Further, Abhyankar and Schuckers (2009) have proposed a wavelet-based approach for the detection of PAs using the fingerprint’s perspiration property. They have used a custom-made fingerprint database to evaluate the performance of their proposed method. The count of pores within a live fingerprint and its counterfeits crafted from various materials can exhibit variability as it is difficult to reflect pores in the spoofs fingerprints. Espinoza and Champod (2011) utilized this dissimilarity as a distinguishing feature to detect PAs on a custom-made dataset. Marcialis et al. (2010) also proposed a similar method that detects pores from multiple fingerprints captured at an interval of five seconds. The number of pores in the subsequent samples is further utilized to detect PAs on a their custom-made database having 8,960 live and 7,760 spoof samples. Manivanan et al. (2010) utilized high-pass filtering followed by correlation filtering techniques Memon et al. (2011) to identify the sweat pores present on fingerprint ridges. Their results demonstrate distinguishable correlation peaks corresponding to the sweat pores on the ridges. Regardless of how the perspiration pattern is employed to detect PAs, its presence is dependent on the temperature of the surrounding environment. Sometimes, even a live finger does not exhibit this property in a dry environment which occasionally prompts the FPAD method to mis-classify this live sample as a spoof. Similarly, the pore-based methods cost high due to the utilization of a high-resolution sensing device that is capable of collecting fingerprint samples with higher dpi (>=>=> =1000 pixels per inch). These factors collectively make perspiration-based and pore-based methods less economical and less convenient to use.
2.2 Handcrafted and statistical feature based-methods
@@
The inherent disparities between the natural attributes of a finger’s skin and counterfeit spoofs, such as variations in color and moisture levels, manifest in the quality of collected samples. This phenomenon prompted researchers to utilize the quality of captured fingerprint samples as a distinguishing feature. This section describes some of the approaches that fall within this category. park et al. (2018) employed a range of statistical features, encompassing deviation, hyper-flatness, average brightness, standard deviation, skewness, variance, hyper-skewness, kurtosis, and differential image features, enabling detection of live and spoof samples using Support Vector Machine (SVM) classifier. The efficacy of the proposed approach is evaluated through experimentation on the ATVSFFp database, comprising 272 genuine and 270 spoof fingerprint samples. In a similar attempt, Xia et al. (2020) developed an image descriptor that captures both gradient characteristics and intensity variance, resulting in a composite feature vector that is fed to an SVM classifier to identify spoofs and live fingerprints The proposed method is validated on LivDet 2011, 2013, and 2015 datasets. Kim (2017) have proposed another image descriptor based on the local coherence property of fingerprints to train an SVM classifier on LivDet 2009, 2011, 2013, and 2015 and ATVSFFp databases. Yuan et al. (2019) introduced a novel approach that utilizes gradient properties for image analysis. This method involves generating image gradient matrices through diverse quantization operators which are derived using the Laplacian operator. These computed matrices are then employed to train a back-propagation neural network using the LivDet 2013 database. Gragnaniello et al. (2013) employed the Weber Local Descriptor (WLD) to extract gradient information and digital excitation from fingerprint samples which are subsequently used to train an SVM classifier. They have evaluated their method using the LivDet 2009 and 2011 databases. Gragnaniello et al. (2015) proposed a novel image descriptor named Local Contrast Phase Descriptor (LCPD) which is a spatial domain and rotation invariant version of Local Phase Quantization (LPQ). The proposed method is validated on LivDet 2011 database. Rattani and Ross (2014) proposed an incremental learning based approach that consists of a novel material detector along with the classifier. This detector divides the input fingerprint samples into three categories: live, spoof, and unknown. Samples identified as ’novel’ are then utilized to train the classifier and evaluated on LivDet 2011 database. In continuation to this work, Rattani et al. (2015) proposed a novel approach that employs a Weibull-calibrated SVM for the classification of fingerprint samples. This approach integrates one-class and binary SVMs, demonstrating a significant improvement over their previous method Rattani and Ross (2014). Since the finger skin and spoof fabrication materials exhibit different levels of elasticity the resultant fingerprints have some differences in terms of quality. The difference in the live and spoof samples fabricated with different materials results in uneven ridge and valley widths and the fabrication process introduces other quality irregularities in the spoofs. In an attempt to develop a quality-based FPAD model, Sharma and Dey (2019) utilized some fingerprint quality features for the detection of PAs. They extracted Ridge and Valley Smoothness (RVS), Ridge and Valley Clarity (RVC), Orientation Certainty Level (OCL), and Frequency Domain Analysis (FDA) to constitute a feature vector for training of Random-Forest (RF) classifier. They have carried out experimental evaluations on LivDet 2009, 2011, 2013, and 2015 databases. Similarly, Sharma and Selwal (2021) proposed a Local Adaptive Binary Pattern (LABP) image descriptor which is a novel variant of the Local Binary Pattern (LBP). The LABP features are combined with the existing Complete Local Binary Pattern (CLBP) and Binary Statistical Image Features (BSIF) to train both an SVM classifier and a Deep Neural Network (DNN). The efficacy of the proposed method is validated across multiple databases, including LivDet 2009, 2011, 2013, and 2015. Rubab Mehboob (2023) proposed a novel method that focuses on liveness features in input fingerprint samples. In this method, the input sample after pre-processing undergoes extraction of two-channel information. The first channel consists of ridge contour in the frequency domain while the second channel consists of ridge contour in the spatial domain. It also extracts minutiae information along with liveness features using the proposed Comprehensive Local Phase Quantization (CLPQ) image descriptor. All the features are used to train an SVM. The proposed method is validated on LivDet 2011, 2013, and 2015 databases. The quality of the fingerprint samples plays a vital role in PA detection by these methods. Some of the methods including Sharma and Dey (2019), Sharma and Selwal (2021) Kim (2017) exhibit good performance while being evaluated in the known material scenario.
@@ 2.3 Deep learning-based-methods
Deep Convolutional Neural Networks (CNNs) excel at extracting minute features from input image samples. These models have shown remarkable classification performance on image datasets such as MNIST Deng (2012), CIFAR Wu et al. (2020), and ImageNet Deng et al. (2009). This ability of the CNNs has prompted researchers to explore their application in PAD. Park et al. (2016) extracted multiple patches from segmented fingerprints and trained a custom-made CNN composed of three convolutional and one fully connected layer. Their method utilizes a voting strategy instead of score-level fusion to compute the overall liveness score of fingerprint samples in LivDet 2009 database. In a similar attempt, Toosi. et al. (2017) applied fingerprint foreground segmentation on fingerprint images followed by patch extraction. The extracted patches are further used to train a pre-trained Alexnet Krizhevsky et al. (2012) CNN classifier using LivDet databases. Frassetto Nogueira et al. (2014) trained two separate SVMs using the features extracted with Local Binary Pattern (LBP) image descriptor and a single-layer CNN respectively. They also incorporated various data augmentation techniques including image rotation and scaling. The limitation of this work is the feature extraction and the classification tasks are performed as separate tasks which do not allow the entire system to learn their parameters together. In another work, Nogueira et al. (2016) utilized pre-trained CNN (VGG, Alexnet, and CNN) along with an SVC classifier for handcrafted features and applied score-level fusion for the detection of PAs on LivDet databases. Uliyan et al. (2020) proposed a method that utilizes a Deep Boltzmann Machine (DBM) for feature extraction and a Restricted Boltzmann Machine (RBM) to establish relationships between these features. Yuan et al. (2022) have presented a multi-modal approach utilizing multiple pre-trained CNNs to detect the PAs on LivDet 2011, 2013, 2015, and NUAA face databases. Chugh and Jain (2021) developed a novel method that crops the patches centered at minutiae points for training the MobileNet classifier. The average liveness score of all the patches is computed to calculate the global liveness score for an input sample. They have tested their method using Michigan State University’s (MSU) FPAD database, and LivDet 2011, 2013, 2015. Zhang et al. (2019) proposed a novel CNN architecture with a series of improved residual connected blocks. It reduces the problem of over-fitting and also takes less processing time. González-Soler et al. (2021) proposed a technique for FPAD in the cross-material scenario that leverages three representations that integrate both global and local information: Bag-of-Words (BoW), Vector Locally Aggregated Descriptors (VLAD), and Fisher Vector (FV). The dense-SIFT descriptor is computed at multiple scales, and the extracted features are encoded using these three feature encoding methods. The encoded features are then utilized to train an SVM classifier. The proposed method is validated on LivDet 2011, 2013, 2015, 2017, and 2019 databases with the FV encoding scheme demonstrating superior performance compared to others. Jing Li (2024) proposed an FPAD method that utilizes the Fisher Vector (FV) method for representing images by aggregating local image features into a single descriptor. The proposed method first transforms the image into local and global feature domains and then extracts discriminating features from them. The local frequency features, local spatial features, and global frequency features are then utilized to classify between live and spoof samples. The proposed method is validated on LivDet 2011, 2013, and 2015 databases. Since, novel fabrication materials are being discovered nowadays, generalizing an open-set FPAD model is a challenging issue. Chugh et al. (2018), in continuation to their previous work Chugh and Jain (2021), proposed an approach for detecting spoofs fabricated with unknown materials. They augmented the fingerprint databases with synthesized fingerprint patches representing the unknown materials spoofs for training the MobileNet classifier. The proposed method is validated on LivDet 2017, ATVSFFp, and MSU-FPAD databases. Generative Adversarial Networks can also be utilized to augment the training database with realistic spoof samples. In a similar attempt, Rai et al. (2024) utilized Wasserstein GAN to generate fingerprint patches to train the DenseNet classifier against the spoofs fabricated with unknown materials. The proposed method is validated using LivDet 2015, 2017, and 2019 databases. In another work, Rai et al. (2025) proposed a novel ensemble of MobileNet and SVC for the detection of PAs in intra-sensor and cross-sensor scenarios of FPAD. The proposed method utilizes MobileNet’s convolutional base for feature extraction and SVC as a classifier. The proposed method is validated on LivDet 2011, 2013, 2015, 2017, and 2019 databases.
Although the Deep learning-based methods are quite effective for image classification problems still, they suffer from low classification accuracy in the area of FPAD while being tested in the cross-material protocol of FPAD. The methods presented in this sub-section prove that the deep learning-based methods are superior as compared with the methods of other categories but their performance is required to be improved in cross-material scenario of FPAD. This objective can be attained by developing a hybrid model that utilizes the power of both convolutional and handcrafted features. The proposed model should also be efficient enough to classify the input sample in a minimum amount of time to be incorporated with the AFRS.